OpenStudy (anonymous):

A falling stone is at a certain instant 128 feet above the ground. 2 seconds later it is only 16 feet above the ground. From what height was it dropped?

OpenStudy (anonymous):

can someone please help me?

OpenStudy (anonymous):

us this Newtonian equation: \[\Delta y = V_0 t + \frac{ 1 }{ 2 }at^2\]

OpenStudy (anonymous):

a = gravity = -32 ft/s^2

OpenStudy (anonymous):

solving it on paper atm

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the answer choices are 139 138 137 140 and 135

OpenStudy (anonymous):

\[\Delta y = y_2 - y_1\] since the stone is dropped, the initial velocity = 0. here are the equations: \[128 - y_0 = 0 - 16t^2\]\[16-y_0 = 0 - 16(t+2)^2\] equate them both for y_0 and solve for t. then you an solve for y_0 by plugging t in. really sorry i gotta go. i hope you can solve from this ^_^ o

OpenStudy (anonymous):

but waht is y0?

OpenStudy (radar):

That is the initial height, what you are trying to solve. Do as @Euler271 posted and you will come up with one of the choices..........Its there just do it.

OpenStudy (radar):